### v2

@@ B.8.6 Checklist-mapped Reference
Consistent with the approach in T1, we generated checklist-mapped references for both human and model outputs. Because T8 has only five checklist items, a single model output extraction prompt was sufficient. The item-level extraction prompt in Tab. 32 and the model output extraction prompt are available in our public GitHub repository. The checklist-mapped reference for the example in Tab. 30 is shown in Tab. 33.
@@ B.9.3 Data Acquisition and Preprocessing
The DiReCT dataset consists of 511 de-identified discharge summaries [78]. From each note, specific sections relevant to the diagnostic process were extracted: chief complaint, history of present illness, past medical history, family history, physical exam, and pertinent results. To ensure high quality data, reasoning annotations were performed by nine licensed clinical physicians and verified for accuracy and completeness by three senior medical experts. Each diagnosis is one of 25 disease categories across five high-level clinical domains.
DiReCT originally contained information that our task has filtered out. The human reference is structured as follows: {o: [z, r, d]}. o is the extracted observation from raw text, z is the rationale to explain why an observation can be related to a diagnosis d, r is the section (from one of input1-6) of the clinical note where o is extracted, and finally, d is the name of the diagnosis. To calculate the #Rubric items of the human reference in Tab. 1 of the main paper, we concatenated this dict into a string and found its token length.
We then selected a smaller, high-quality, and representative subset from the dataset with the criteria of diversity and difficulty described in §3.2. To maintain diversity, we aimed to sample equally from all 25 final diagnosis. This promotes a balanced distribution that supports generalizable evaluation across multiple clinical domains. We first selected six samples per diagnosis type, resulting in an initial pool of 150 samples. Next, we prioritized cases with longer and more complex reasoning chains, as these are more likely to challenge model capabilities in clinical inference. From the 150 samples, we identified 100 with the longest human-generated reference outputs, reflecting the depth of clinical reasoning involved. Finally, we manually verified that each sample met both diversity and difficulty thresholds. Tab. 34 shows the diagnosis distribution of our final 100 samples across 18 diagnoses.
@@ B.9.4 Illustrative Example
Due to the proprietary nature of the data, we cannot provide a specific example. However, Tab. 35 displays the model prompt. The model is instructed to diagnose the patient and provide supporting evidence in a structured dict_reasoning format. To ensure alignment with human references and enable verification, we constrained the model to select from a predefined list of possible diagnoses. This was necessary because, during early testing, the model often generated overly specific diagnoses— "Hypertension in the setting of atrial fibrillation" instead of "Hypertension"—which could not be matched against the simpler human reference.
@@ B.10.4 Illustrative Example
Due to the proprietary nature of the data, we cannot provide a specific example. However, in Tab. 36, we provide the model prompt used in obtaining model outputs.
@@
You are an ESG analyst. You will read the ESG report and sustainability report of a company. Based on the reports, write a summary of the company’s ESG performance. The summary should focus on the company’s performance on key ESG issues, and should include key details (e.g., strategies and initiatives). Word limit: 300 words.
@@ B.10.5 Evaluation Rubric
The ESG report summaries discuss the company’s performance on ESG key issues. While the ESG report of a company details its performance on ESG key issues that are important to the corresponding sector, a well-written summary should cover the most significant key issues relevant to the company and give a precise overview of the company’s performance on these issues. We build the evaluation rubrics based on MSCI’s ESG scoring framework, which includes 33 key issues across 10 themes. [^29] https://www.msci.com/documents/1296102/34424357/MSCI+ESG+Ratings+Methodology.pdf The discussion of key issues is sector-specific, [^30] https://www.msci.com/our-solutions/esg-investing/esg-industry-materiality-map and therefore in this task we have different rubrics for different data samples, as shown in Tab. 37.
@@ B.10.6 Checklist-mapped Reference
For each sample, to map the human-written summary into the reference checklist, we prompt the checklist mapper with the instruction that contains all the rubric items and their descriptions for the sector the company belongs to. The items are not separately extracted, as it is easier for the checklist mapper to distinguish between items that originate from the same key issue when they are presented together. We show the prompt for checklist mapping in LABEL:tab:T11-checklist-mappedReference-prompt.

### v3

@@ B.8.6 Checklist-mapped Reference
Consistent with the approach in T1, we generated checklist-mapped references for both human and model outputs. Because T8 has only five checklist items, a single model output extraction prompt was sufficient. The item-level extraction prompt in Tab.˜32 and the model output extraction prompt will be available in our public GitHub repository. The checklist-mapped reference for the example in Tab.˜30 is shown in Tab.˜33.
@@ B.9.3 Data Acquisition and Preprocessing
The DiReCT dataset consists of 511 de-identified discharge summaries Wang et al. [2024]. From each note, specific sections relevant to the diagnostic process were extracted: chief complaint, history of present illness, past medical history, family history, physical exam, and pertinent results. To ensure high quality data, reasoning annotations were performed by nine licensed clinical physicians and verified for accuracy and completeness by three senior medical experts. Each diagnosis is one of 25 disease categories across five high-level clinical domains.
DiReCT originally contained information that our task has filtered out. The human reference is structured as follows: {o: [z, r, d]}. o is the extracted observation from raw text, z is the rationale to explain why an observation can be related to a diagnosis d, r is the section (from one of input1-6) of the clinical note where o is extracted, and finally, d is the name of the diagnosis. To calculate the #Rubric items of the human reference in Tab.˜1 of the main paper, we concatenated this dict into a string and found its token length.
We then selected a smaller, high-quality, and representative subset from the dataset with the criteria of diversity and difficulty described in §3. To maintain diversity, we aimed to sample equally from all 25 final diagnosis. This promotes a balanced distribution that supports generalizable evaluation across multiple clinical domains. We first selected six samples per diagnosis type, resulting in an initial pool of 150 samples. Next, we prioritized cases with longer and more complex reasoning chains, as these are more likely to challenge model capabilities in clinical inference. From the 150 samples, we identified 100 with the longest human-generated reference outputs, reflecting the depth of clinical reasoning involved. Finally, we manually verified that each sample met both diversity and difficulty thresholds. Tab.˜34 shows the diagnosis distribution of our final 100 samples across 18 diagnoses.
@@ B.9.4 Illustrative Example
Due to the proprietary nature of the data, we cannot provide a specific example. However, Tab.˜35 displays the model prompt. The model is instructed to diagnose the patient and provide supporting evidence in a structured dict_reasoning format. To ensure alignment with human references and enable verification, we constrained the model to select from a predefined list of possible diagnoses. This was necessary because, during early testing, the model often generated overly specific diagnoses— "Hypertension in the setting of atrial fibrillation" instead of "Hypertension"—which could not be matched against the simpler human reference.
@@ B.10.4 Illustrative Example
Due to the proprietary nature of the data, we cannot provide a specific example. However, in Tab.˜36, we provide the model prompt used in obtaining model outputs.
@@
{internallinenumbers*} You are an ESG analyst. You will read the ESG report and sustainability report of a company. Based on the reports, write a summary of the company’s ESG performance. The summary should focus on the company’s performance on key ESG issues, and should include key details (e.g., strategies and initiatives). Word limit: 300 words.
@@ B.10.5 Evaluation Rubric
The ESG report summaries discuss the company’s performance on ESG key issues. While the ESG report of a company details its performance on ESG key issues that are important to the corresponding sector, a well-written summary should cover the most significant key issues relevant to the company and give a precise overview of the company’s performance on these issues. We build the evaluation rubrics based on MSCI’s ESG scoring framework, which includes 33 key issues across 10 themes. [^29] https://www.msci.com/documents/1296102/34424357/MSCI+ESG+Ratings+Methodology.pdf The discussion of key issues is sector-specific, [^30] https://www.msci.com/our-solutions/esg-investing/esg-industry-materiality-map and therefore in this task we have different rubrics for different data samples, as shown in Tab.˜37.
@@ B.10.6 Checklist-mapped Reference
For each sample, to map the human-written summary into the reference checklist, we prompt the checklist mapper with the instruction that contains all the rubric items and their descriptions for the sector the company belongs to. The items are not separately extracted, as it is easier for the checklist mapper to distinguish between items that originate from the same key issue when they are presented together. We show the prompt for checklist mapping in Tab.˜38.